OpenStudy (anonymous):

A student scores 71 on a test. The maximum points is 100. what score does the student need to maintain at least an average of 80? Just tell me how to solve the question.

OpenStudy (jdoe0001):

something tells me the info is incomplete :/ and average value, assumes there's a serie of accumulated values that if added together and divided by some relevant divisor, will yield an "average" thus far we only have 1 value, 1 test

OpenStudy (jdoe0001):

hmmm, I see, he scores 71 on a prior test, no the 100% one, ok hhehe

OpenStudy (anonymous):

average of test scores = sum of test scores divided by the number of tests

OpenStudy (whpalmer4):

If we have 2 tests, each worth 100 points, the 1st test he scores 71, and he wants the average to be 80, then \[\frac{\frac{71}{100}+\frac{x}{100} }{2}= \frac{80}{100}\]

OpenStudy (anonymous):

This is the question...sorry, computer froze on me. A student scores 71 out of 100 on a test. if the maximum score on the next test is also 100 points what score does the student need to maintain at least an average of 80. it seems that I overlooked the next test part....

OpenStudy (whpalmer4):

In other words, to have an average of 80%, there are a total of 200 possible points, and he needs to accumulate 80%*200 = 160 points between the two tests. He's already got 71, so 160-71 is the score he needs.

OpenStudy (whpalmer4):

Yet another way of thinking about it: if the tests are all worth the same amount, you need to offset the number of points below the desired average on the tests taken so far with points above the desired average on tests taken in the future. For example, you've got 5 tests each worth 100 points, you want to have an average of 85%, and so far you've gotten 78, 81, and 72 points on the 3 tests you've taken. That's -2 + 1 - 8 = -9 points you need to make up on the next two tests to get to 80%. If you score 80+80+9 = 169 points over the two remaining tests, you'll have 78+81+72+80+80+9 = 400 points, or 80% of 500 points.

OpenStudy (anonymous):

Okay...you provided me with example.